# Equivariant CM minimization for extremal manifolds

### v1

@@ 2.1.2. Equivariant and twisted models
We introduce a subclass of models that are compatible with vertical torus actions on $(X,L)\to\operatorname{Spec}K$. Let $T$ be a $K$-torus in $\mathrm{Aut}_{K}(X,L)$, meaning $T\cong\mathbb{G}_{m,K}^{d}$ for some integer $d\geq 0$.
@@
Choose a uniformizer $t$ in the discrete valuation ring, that is, an element in $R$ such that its valuation $v(t)$ generates the valuation group $v(K^{*})$ in $\mathbb{R}$; (in the case $R=\mathbb{C}[\![t]\!]$, the element $t$ itself is a uniformizer). Let $(\mathcal{X},\mathcal{L})$ be a model of $(X,L)$ and let $\beta$ be a vertical $\mathbb{G}_{m,K}$-action on $(X,L)$. The choice of uniformizer $t$ gives rise to an automorphism $\beta(t):(X,L)\to(X,L)$.
@@
($\mathcal{F}^{\lambda}_{(Y\times\mathbb{C},H)_{\beta}}S_{k}=\bigoplus_{\mu\geq\lambda}(S_{k})_{\mu},$
@@
($\mathcal{F}^{\lambda}_{(r)}S_{kr}=\{\sigma\;|\;\sigma\cdot t^{-\lambda}\in\mathcal{S}_{kr}(\mathcal{F})\},$
@@
where $P_{k}=\{\nu_{\sigma}(f)\;|\;f\in S_{k}\}\subset\mathbb{Z}^{n}$ and $\nu_{\sigma}(f)$ is the multi-degree of the lowest-order term of the power series $q$ in $\mathbb{Z}[\![z_{1},\dots,z_{n}]\!]$ such that $f=\sigma^{k}q$ on $U$ with respect to the lexicographic order. We call $P$ the Okounkov body of $(Y,H)$ centered around $p$, induced by the data $(U,z,\sigma)$. Important properties are $|P_{k}|=\dim S_{k}$ and $\mathrm{vol}(P)=n!\mathrm{vol}(L)$. From [Szé15, Lemma 6], if follows that for sufficiently large integers $k$ and $r$, we have $\Delta_{k-1}\cap\mathbb{Z}^{n}\subset P_{kr}$, where $\Delta_{l}=\{(x_{1},\dots,x_{n})\in\mathbb{R}^{n}\;|\;x_{i}\geq 0\text{ and }\sum_{i=1}^{n}x_{i}\leq l\}$ denotes the closed solid $n$-simplex of side length $l>0$ In particular, we have $\Delta_{1/r}\subset P$.
@@
$g_{\mathcal{F},k}(a)$ $=\max\{\lambda\in\mathbb{Z}\;|\;a\in P_{k}(\tfrac{\lambda}{k},\mathcal{F})\},$
@@
The function $G_{\mathcal{F}}$ is concave and upper semi-continuous on $P$ and is finite and continuous on the interior of $P$. A key property is that one may express the leading coefficient of the $p$th-order weight function of $\mathcal{F}$ as
@@
$||\mathcal{F}||^{2}=\frac{1}{\mathrm{vol}(P)}\int_{P}(G_{\mathcal{F}}-\overline{G}_{\mathcal{F}})^{2},$
@@ A $\beta$-invariant filtration $\mathcal{F}$ gives rise to weight-space decompositions
$||\mathcal{F}_{\beta}||^{2}=||\mathcal{F}||^{2}+2\langle\mathcal{F},\beta\rangle+||\beta||^{2}.$
@@
A modification of [Nys12, Lemma 3.3] and [BC11, Theorem A] gives rise to the following relation and asymptotics between the weight pieces, lattice functions and concave transforms:
@@
$\sum_{\lambda,\mu\in\mathbb{Z}}T(\lambda,\mu)\dim((\mathcal{F}^{\lambda}S_{k}/\mathcal{F}^{\lambda+1}S_{k})_{\mu})=\sum_{a\in P_{k}}T(g_{\mathcal{F},k}(a),g_{\beta,k}(a)).$
@@
$\sum_{\lambda,\mu\in\mathbb{Z}}$ $T(\lambda,\mu)\dim((\mathcal{F}^{\lambda}S_{k}/\mathcal{F}^{\lambda+1}S_{k})_{\mu}),$
@@
The subset $P(t,\mathcal{F})$ in $P$ is the Okounkov subbody of the graded subalgebra $S(t,\mathcal{F})=\bigoplus_{k\geq 0}S_{k}(t,\mathcal{F})$ for $S_{k}(t,\mathcal{F})=\mathcal{F}^{\lceil tk\rceil}S_{k}$. Hence, by [BC11, Equation 1.8] and [LM09, Theorem 2.13], in the case where $T$ is smooth, we have
@@
$=\lim_{k\to\infty}\frac{\sum_{a\in P_{k}}T\big{(}\frac{g_{\mathcal{F},k}(a)}{k},\frac{g_{\beta,k}(a)}{k}\big{)}}{k^{n}}.$
where we may change the limit with the integral as the volume of $P(t,\mathcal{F})$ is bounded uniformly by the volume of $P$. The result for general continuous functions follows by density. ∎
@@ 2.2.3. Torus-invariant filtrations
A filtration $\mathcal{F}$ is $T$-invariant if and only if it is $\beta$-invariant for all $\beta$ in $N_{\mathbb{Z}}(T)$. The mixed-square weight $\langle\mathcal{F},\beta\rangle$ extends bilinearly to the real vector space $N_{\mathbb{R}}(T)=N_{\mathbb{Z}}(T)\otimes\mathbb{R}$. This allows us to make the following definition:
@@
$||\mathcal{F}||_{T}^{2}=\inf_{\xi\in N_{\mathbb{R}}(T)}(||\mathcal{F}||^{2}+2\langle\mathcal{F},\xi\rangle+||\xi||^{2}),$
where $||\xi||^{2}=\langle\xi,\xi\rangle$ is the natural extension of the norm to $N_{\mathbb{R}}(T)$.
@@
$||\mathcal{F}||_{T}^{2}=||\mathcal{F}||^{2}-2\langle\mathcal{F},\xi(\mathcal{F})\rangle+||\xi(\mathcal{F})||^{2},$
@@
$||\xi(\mathcal{F})+\zeta||^{2}-2\langle\mathcal{F},\xi(\mathcal{F})+\zeta\rangle=||\xi(\mathcal{F})||^{2}+||\zeta||^{2}-2\langle\mathcal{F},\xi(\mathcal{F})\rangle.$
This is bounded from below by $||\xi(\mathcal{F})||^{2}-2\langle\mathcal{F},\xi(\mathcal{F})\rangle$, as desired. ∎
@@ 2.2.4. Good filtrations
An issue with non-finitely-generated filtrations is that, in general, we have no control over subleading coefficients of the weight function. This makes it difficult to define the polynomial Donaldson–Futaki invariant of general filtrations, so we introduce a subclass of filtrations for which we have control over the weight function as a polynomial.
@@
A filtration $\mathcal{F}$ of $(Y,H)$ is called good if there is a polynomial $b(t)$ over $\mathbb{Q}$ such that $w_{\mathcal{F}}(k)=b(k)$ for sufficiently large integers $k$. The Donaldson–Futaki invariant of a good filtration $\mathcal{F}$ is defined as
@@
where $\mathrm{Fut}(\beta)=\mathrm{DF}(\beta)$ is the Futaki invariant of $\beta$ [Fut83].
Remark 2.23 ([Hat24, Remark 2.21]).
The definition of the Donaldson-Futaki invariant of a good filtration in principle differs from the definition in [Szé15, Definition 3] in the sense that
$\mathrm{DF}(\mathcal{F})\leq\liminf_{r\to+\infty}\mathrm{DF}(\mathcal{F}_{(r)}),$
for good filtrations $\mathcal{F}$. It is not known whether this is an equality in general; this is a problem about the convergence of subleading coefficients of the weight function, and is related to the non-Archimedean entropy approximation conjecture [BJ18, Conjecture 2.5] and [Li22, Conjecture 1.8].
@@ 3. Relative stability of extremal polarized manifolds
Let $(Y,H)$ be a smooth polarized variety and let $T$ in $\mathrm{Aut}(Y,H)$ be a maximal complex torus of rank $d$. The goal of this section is to prove that if $(Y,H)$ is extremal, then it is relatively K-polystable with respect to good filtrations. This is second step outlined in Section 1.3 used in the proof of 1.2. We start by introducing the relevant numerical invariant.
@@
where the inequality is strict if and only if $||\mathcal{F}||_{T}>0$.
This is a simultaneous generalization of [Hat24, Theorem 2.18], [Szé15, Theorem 10] and [SS11, Theorem 4] allowing automorphisms and filtrations. In the case where $(Y,H)$ admits a constant scalar curvature Kähler (cscK) metric, the Futaki invariant in Lemma 2.22 vanishes, giving rise to the following:
@@
where the inequality is strict if and only if $||\mathcal{F}||_{T}>0$.
@@ 3.1. Székelyhidi’s formula
where $a_{0}$ and $b_{0}$ are as in Equation (2.5). Choose a point $p$ in $Y$ and write $(\widehat{Y},\widehat{H}_{m})=(\mathrm{Bl}_{p}Y,mH-E)$ for the blowup with exceptional divisor $E$, which is a polarized variety for sufficiently large integers $m$. Given a filtration $\mathcal{F}$ of $(Y,H)$, we define a new filtration $\widehat{\mathcal{F}}$ of $(\widehat{Y},\widehat{H}_{m})$ by
(3.1) ^$\widehat{\mathcal{F}}^{\lambda}H^{0}\big{(}\widehat{Y},k\widehat{H}_{m}\big{)}=\{\sigma\;|\;i(\sigma)\in\mathcal{F}^{\lambda}H^{0}(Y,kmH)\},$
@@
From now on, we will assume $n\geq 2$. We will refer to the case $n\leq 1$ in Remark 3.17 when discussing the stability results. Székelyhidi’s formula is the following asymptotic bound on the Chow weight on the blowup:
Theorem 3.4 ([Szé15, Page 18]).
Let $P$ be an Okounkov body of $(Y,H)$ centered at $p$. Then
$\mathrm{Chow}_{r}(\widehat{\mathcal{F}})\leq\mathrm{Chow}_{mr}(\mathcal{F})+\Big{(}\nu_{m}(p,\mathcal{F})-\frac{b_{0}}{a_{0}}\Big{)}\frac{m^{1-n}}{2(n-2)!a_{0}}+o(m^{-n}r^{0}),$
as $r\to+\infty$ and subsequently $m\to+\infty$, where
(3.2) $\nu_{m}(p,\mathcal{F})=\frac{(n-2)!a_{0}}{m^{1-n}}\bigg{(}\int_{\partial_{0}\Delta_{\frac{1}{m}}}G_{\mathcal{F}}-\frac{1}{\sqrt{n}}\int_{\partial_{1}\Delta_{\frac{1}{m}}}G_{\mathcal{F}}\bigg{)}.$
Here, $\partial_{0}\Delta_{l}$ denotes the union of boundary faces meeting in the origin the solid closed $n$-simplex $\Delta_{l}\subset\mathbb{R}^{n}$ of length $l$ and $\partial_{1}\Delta_{l}$ denotes the remaining face. The value in Equation (3.2) depends on the choice of Okounkov body; however, we always have $\nu_{m}(\mathcal{F},p)=O(m^{0})$ as $m\to+\infty$.
@@
We next establish the a relative blowup formula, extending Theorem 3.4, needed in the proof of Theorem 3.16. Given a $\mathbb{C}^{*}$-action $\beta$ on $(Y,H)$ and a $\beta$-invariant filtration $\mathcal{F}$ be $(Y,H)$, we define
@@
^$\langle\widehat{\mathcal{F}},\widehat{\beta}\rangle_{r}=\langle\mathcal{F},\beta\rangle_{mr}m^{2}+o(m^{2-n}r^{0}).$
@@
By asymptotic Riemann-Roch we have
@@
Given a choice of Okounkov body $P$ for $(Y,H)$ centered at $p$ constructed out of data $(U,z,\sigma)$, one obtains an Okounkov body
$\widehat{P}=\overline{mP\setminus\Delta_{1}}$
for $(\widehat{Y},\widehat{H}_{m})$ constructed out of lifted data $(\widehat{U},\widehat{z},\widehat{\sigma^{m}})$. A direct computation shows that $G_{\widehat{\mathcal{F}}}(x)=mG_{\mathcal{F}}(\frac{x}{m})$ for all $x$ in $\widehat{P}$. Using this and Equation (2.7), we have
@@
^$\langle\widehat{\mathcal{F}},\widehat{\beta}\rangle=\langle\mathcal{F},\beta\rangle m^{2}+O(m^{2-n})$
as $m\to+\infty$, which matches the formula proven by Stoppa–Szkékelyhidi [SS11, Page 8] for test configurations.
@@
where $(\beta_{i})$ is any choice of $\mathbb{Z}$-basis of $N_{\mathbb{Z}}(T)$. In the case of test configurations, Equation (3.4) captures the weight relevant to relative Chow stability.
@@
This is an equality whenever $\mathcal{F}$ is finitely generated as the approximation $\mathcal{F}_{(r)}$ then is $\mathcal{F}$, viewed as a filtration of $(X,rL)$. Moreover, as ($\mathcal{F}^{\lambda}_{(r)}S_{r}=\mathcal{F}^{\lambda}S_{r}$ for any filtration $\mathcal{F}$, we have $\langle\mathcal{F}_{(r)},\beta_{i}\rangle_{1}=\langle\mathcal{F},\beta_{i}\rangle_{r}$, proving the result. ∎
Assuming now that $p$ is fixed by $T$, we obtain a torus $\widehat{T}$ in $\mathrm{Aut}(\widehat{Y},\widehat{H}_{m})$ which leaves $\widehat{\mathcal{F}}$ invariant. Combining Theorem 3.4 and Proposition 3.5, we obtain the following relative formula:
Proposition 3.8.
We have
$\mathrm{Chow}_{r,T}(\widehat{\mathcal{F}})\leq\mathrm{Chow}_{mr,T}(\mathcal{F})+\Big{(}\nu_{m,T}(p,\mathcal{F})-\frac{b_{0}}{a_{0}}\Big{)}\frac{m^{1-n}}{2(n-2)!a_{0}}+o(m^{-n}r^{0})$
@@
(3.5) $\nu_{m,T}(p,\mathcal{F})=\nu_{m}(p,\mathcal{F})-\sum_{i=1}^{d}\frac{\langle\mathcal{F},\beta_{i}\rangle}{\langle\beta_{i},\beta_{i}\rangle}\Big{(}\nu_{m}(p,\beta_{i})-\frac{b_{i,0}}{a_{0}}\Big{)},$
for any $\mathbb{Z}$-basis $(\beta_{i})$ of $N_{\mathbb{Z}}(T)$ and $w_{\beta_{i}}(k)=b_{i,0}k^{n+1}+O(k^{n})$.
@@
In order to use the formula in Proposition 3.8 for an instability argument on the blowup, we need to find an invariant point $p$ for which the relative weight in Equation (3.5) is small for large enough $m$. For this, we establish a $T$-invariant version of results of Boucksom [Szé15, Section 8].
Let $S^{\prime}$ be a graded subalgebra of the ring of sections $S=\bigoplus_{k\geq 0}H^{0}(Y,kH)$ not equal to $\mathbb{C}$. Recall the volume of $S^{\prime}$ is the limit
@@
where $\mathfrak{bs}(S_{k}^{\prime})$ is the base-ideal of the linear series $S_{k}^{\prime}$ in $S$. The proofs of Lemmas 3.11 and 3.12 follow verbatim the proofs of [Szé15, Lemma 22] and [Szé15, Lemma 21] noting that the divisorial valuations and their centers constructed in the proofs are invariant, so we omit the details.
@@
Let $\Lambda<e_{\mathrm{max}}(\mathcal{F})$ be a real number. There exists a $T$-invariant point $p$ in $X$ such that for any Okounkov body $P$ centered at $p$, there is an integer $m_{1}\geq 1$ such that
$G_{\mathcal{F}}\leq\Lambda\quad\text{on}\quad\Delta_{\frac{1}{m_{1}}}\subset P.$

### v2

@@ 2.1.2. Equivariant and twisted models
We introduce a subclass of models that are compatible with vertical torus actions on $(X,L)\to\operatorname{Spec}K$. Let $T$ be a $K$-torus in $\mathrm{Aut}_{K}(X,L)$, meaning that $T\cong\mathbb{G}_{m,K}^{d}$ for some integer $d\geq 0$.
@@
Choose a uniformizer $t$ in the discrete valuation ring, that is, an element in $R$ such that its valuation $\mathsf{v}(t)$ generates the valuation group $\mathsf{v}(K^{*})$ in $\mathbb{R}$; (in the case $R=\mathbb{C}[\![t]\!]$, the element $t$ itself is a uniformizer). Let $(\mathcal{X},\mathcal{L})$ be a model of $(X,L)$ and let $\beta$ be a vertical $\mathbb{G}_{m,K}$-action on $(X,L)$. The choice of uniformizer $t$ gives rise to an automorphism $\beta(t):(X,L)\to(X,L)$.
@@
Remark 2.9.
In the more geometrical case of families $(X,L)\xrightarrow{\pi}\Delta^{*}$ over the punctured disk in $\mathbb{C}$, one would choose a vertical $\mathbb{C}^{*}$-action $\beta$ on $(X,L)$, giving rise to an automorphism $(X,L)$, sending a point $(x,\xi)\mapsto\beta(\pi(x))(x,\xi)$ for $x$ in $X$ and $\xi$ in $L_{x}$. The twist of a filling $(\mathcal{X},\mathcal{L})\to\Delta$ to the whole disc would be defined in the same way by composing the identification with the automorphism.
@@
$\mathcal{F}^{\lambda}_{(Y\times\mathbb{C},H)_{\beta}}S_{k}=\bigoplus_{\mu\geq\lambda}(S_{k})_{\mu},$
@@
$\mathcal{F}^{\lambda}_{(r)}S_{kr}=\{\sigma\;|\;\sigma\cdot t^{-\lambda}\in\mathcal{S}_{kr}(\mathcal{F})\},$
@@
where $P_{k}=\{\nu_{\sigma}(f)\;|\;f\in S_{k}\}\subset\mathbb{Z}^{n}$ and $\nu_{\sigma}(f)$ is the multi-degree of the lowest-order term of the power series $q$ in $\mathbb{Z}[\![z_{1},\dots,z_{n}]\!]$ such that $f=\sigma^{k}q$ on $U$ with respect to the lexicographic order. We call $P$ the Okounkov body of $(Y,H)$ centered around $p$, induced by the data $(U,z,\sigma)$. Important properties are $|P_{k}|=\dim S_{k}$ and $\mathrm{vol}(P)=n!\mathrm{vol}(L)$. From [Szé15, Lemma 6], if follows that for sufficiently large integers $k$ and $r$, we have $\Delta_{k-1}\cap\mathbb{Z}^{n}\subset P_{kr}$, where $\Delta_{l}=\{(x_{1},\dots,x_{n})\in\mathbb{R}^{n}\;|\;x_{i}\geq 0\text{ and }\sum_{i=1}^{n}x_{i}\leq l\}$ denotes the closed solid $n$-simplex of side length $l>0$ . In particular, we have $\Delta_{1/r}\subset P$.
@@
$g_{\mathcal{F},k}(a)$ $=\max\big\{\lambda\in\mathbb{Z}\;\big|\;a\in P_{k}\big(\tfrac{\lambda}{k},\mathcal{F}\big)\big\},$
@@
The function $G_{\mathcal{F}}$ is concave and upper semi-continuous on $P$ and is finite and continuous on the interior of $P$. A key property (see [Nys12, Szé15]) is that one may express the leading coefficient of the $p$th-order weight function of $\mathcal{F}$ as
@@
$\mathcal{F}||^{2}=\frac{1}{\mathrm{vol}(P)}\int_{P}(G_{\mathcal{F}}-\overline{G}_{\mathcal{F}})^{2},$
@@ A $\beta$-invariant filtration $\mathcal{F}$ gives rise to weight-space decompositions
[…]
@@
A modification of [Nys12, Lemma 3.3] and [BC11, Theorem A] gives rise to the following relation and asymptotic formula between the weight pieces, lattice functions and concave transforms:
@@
$\sum_{\lambda,\mu\in\mathbb{Z}}T(\lambda,\mu)\dim\big((\mathcal{F}^{\lambda}S_{k}/\mathcal{F}^{\lambda+1}S_{k})_{\mu}\big)=\sum_{a\in P_{k}}T(g_{\mathcal{F},k}(a),g_{\beta,k}(a)).$
@@
$\sum_{\lambda,\mu\in\mathbb{Z}}$ $T(\lambda,\mu)\dim\big((\mathcal{F}^{\lambda}S_{k}/\mathcal{F}^{\lambda+1}S_{k})_{\mu}\big),$
@@
The subset $P(t,\mathcal{F})$ in $P$ is the Okounkov subbody of the graded subalgebra $S(t,\mathcal{F})=\bigoplus_{k\geq 0}S_{k}(t,\mathcal{F})$ for $S_{k}(t,\mathcal{F})=\mathcal{F}^{\lceil tk\rceil}S_{k}$. In the case where $T$ is continuously differentiable, writing $DT(t,s):\mathbb{R}^{2}\to\mathbb{R}$ for the total derivative of $T$ at a point $(t,s)$, we have
@@
$=\lim_{k\to\infty}\frac{\sum_{a\in P_{k}}T\big(\frac{g_{\mathcal{F},k}(a)}{k},\frac{g_{\beta,k}(a)}{k}\big)}{k^{n}},$
where we used [BC11, Equation 1.8] and [LM09, Theorem 2.13]. Note that we were allowed to change the limit with the integral as the volume of $P(t,\mathcal{F})$ is bounded uniformly by the volume of $P$. The result for general continuous functions follows by a density argument. ∎
@@ 2.2.3. Torus-invariant filtrations
A filtration $\mathcal{F}$ is $T$-invariant if and only if it is $\beta$-invariant for all $\beta$ in $N_{\mathbb{Z}}(T)$. Furthermore, the inner product $\langle\mathcal{F},\beta\rangle$ extends bilinearly to the real vector space $N_{\mathbb{R}}(T)=N_{\mathbb{Z}}(T)\otimes\mathbb{R}$. This allows us to make the following definition:
@@
[…]
where $\xi||^{2}=\langle\xi,\xi\rangle$ is the natural extension of the norm to $N_{\mathbb{R}}(T)$.
@@
[…]
@@
[…]
This is bounded from below by $\xi(\mathcal{F})||^{2}-2\langle\mathcal{F},\xi(\mathcal{F})\rangle$, as desired. ∎
@@ 2.2.4. Good filtrations
An issue with non-finitely-generated filtrations is that, in general, we have no control over subleading coefficients of the weight function. This makes it difficult to define the polynomial Donaldson–Futaki invariant of general filtrations. To fix this, we introduce a subclass of filtrations for which we have control over the weight function as a polynomial.
@@
A filtration $\mathcal{F}$ of $(Y,H)$ is called good if there is a polynomial $b(t)$ over $\mathbb{Q}$ such that $w_{\mathcal{F}}(k)=b(k)$ for sufficiently large integers $k$. The Donaldson–Futaki invariant of a good filtration $\mathcal{F}$ is defined to be
@@
where $\mathrm{Fut}(\beta)=\mathrm{DF}(\beta)$ is the Futaki invariant of $\beta$ [Fut83, Don02].
@@ 3. Relative stability of extremal polarized manifolds
Let $(Y,H)$ be a smooth polarized variety and let $T$ in $\mathrm{Aut}(Y,H)$ be a maximal complex torus of rank $d$. The goal of this section is to prove that if $(Y,H)$ is extremal, then it is relatively K-polystable with respect to good filtrations. This is second step outlined in Section 1.3, and used in the proof of 1.2. We start by introducing the relevant numerical invariant.
@@
where the inequality is strict if and only if $\mathcal{F}||_{T}>0$.
This is a simultaneous generalization of [Hat24, Theorem 2.18], [Szé15, Theorem 10] and [SS11, Theorem 4] allowing automorphisms and filtrations. In the case where $(Y,H)$ admits a constant scalar curvature Kähler (cscK) metric, the Futaki invariant in Lemma 2.23 vanishes (due to [Fut83]), giving rise to the following consequence:
@@
where the inequality is strict if and only if $\mathcal{F}||_{T}>0$.
@@ 3.1. Székelyhidi’s formula
where $a_{0}$ and $b_{0}$ are as in Equation (2.5).
Choose a point $p$ in $Y$ and write $(\widehat{Y},\widehat{H}_{m})=(\mathrm{Bl}_{p}Y,mH-E)$ for the blowup with exceptional divisor $E$, which is a polarized variety for all $m\geq m_{0}$ for a sufficiently large integer $m_{0}$. Given a filtration $\mathcal{F}$ of $(Y,H)$, we define a new filtration $\widehat{\mathcal{F}}$ of $(\widehat{Y},\widehat{H}_{m})$ by
(3.1) $\widehat{\mathcal{F}}^{\lambda}H^{0}\big(\widehat{Y},k\widehat{H}_{m}\big)=\{\sigma\;|\;i(\sigma)\in\mathcal{F}^{\lambda}H^{0}(Y,kmH)\},$
@@
Given a choice of Okounkov body $P$ for $(Y,H)$ centered at $p$, constructed out of data $(U,z,\sigma)$, one obtains an Okounkov body
$\widehat{P}=\overline{mP\setminus\Delta_{1}},$
[…]
From now on, we will assume $n\geq 2$. We will refer to the case $n\leq 1$ in Remark 3.17 when discussing the stability results. Székelyhidi’s formula states that the Chow weight of the filtration $\widehat{\mathcal{F}}$ on the blowup is asymptotically bounded by the Chow weight of the tautological approximation $(\mathcal{F}_{(r)})$ of $\mathcal{F}$ (see Example 2.12) in the following way:
Theorem 3.4 ([Szé15, Lemma 12]).
Let $P$ be an Okounkov body of $(Y,H)$ centered at $p$ and assume that $G_{\mathcal{F}}\leq\Lambda$ on $\Delta_{\frac{1}{m}}$ for a real number $\Lambda$, where $\Delta_{l}$ denotes the solid closed $n$-simplex of length $l$ in $\mathbb{R}^{n}$. Then,
$w_{\widehat{\mathcal{F}}}(r)\leq w_{\mathcal{F}_{(mr)}}(1)-mr^{n}\int_{\Delta_{\frac{1}{m}-\frac{n}{r}}}G_{\mathcal{F}}-\frac{\Lambda(3n-1)}{2(n-1)!}m^{1-n}r^{n-1}+O(m^{0}r^{n-2})$
as $r\to+\infty$ and subsequently $m\to+\infty$. In particular, we have
$\mathrm{Chow}_{r}(\widehat{\mathcal{F}})\leq\mathrm{Chow}_{1}(\mathcal{F}_{(rm)})+\Big(\nu(p,\mathcal{F})-\frac{b_{0}}{a_{0}}\Big)\frac{m^{1-n}}{2(n-2)!a_{0}}+o(m^{-n}r^{0})$
as $r\to+\infty$ and $m\to+\infty$, where
$\nu(p,\mathcal{F})=(n-2)!a_{0}\bigg(\frac{\Lambda(3n-1)}{2(n-1)!}-\frac{r}{m^{1-n}}\int_{\Delta_{\frac{1}{m}}\setminus\Delta_{\frac{1}{m}-\frac{n}{r}}}G_{\mathcal{F}}\bigg).$
In the proof of Theorem 3.16, we will use specific choices of $p$ and $\Lambda$, coming from Corollary 3.14.
@@
We next establish the a relative blowup formula, extending Theorem 3.4, which will be needed in the proof of Theorem 3.16. Given a $\mathbb{C}^{*}$-action $\beta$ on $(Y,H)$ and a $\beta$-invariant filtration $\mathcal{F}$ be $(Y,H)$, we define
@@
$\langle\widehat{\mathcal{F}},\widehat{\beta}\rangle_{r}=\langle\mathcal{F},\beta\rangle_{mr}m^{2}+o(m^{2-n}r^{0}).$
@@
By the asymptotic Riemann-Roch formula we have
@@
Using $G_{\widehat{\mathcal{F}}}(x)=mG_{\mathcal{F}}(\frac{x}{m})$ on $\widehat{P}=\overline{mP\setminus\Delta_{1}}$ and Equation (2.7), we have
@@
$\langle\widehat{\mathcal{F}},\widehat{\beta}\rangle=\langle\mathcal{F},\beta\rangle m^{2}+O(m^{2-n})$
as $m\to+\infty$. This matches the formula proven by Stoppa–Szkékelyhidi [SS11, Page 8] for test configurations.
@@
where $(\beta_{i})$ is any choice of $\mathbb{Z}$-basis of $N_{\mathbb{Z}}(T)$. In the case of test configurations, Equation (3.3) captures the weight relevant for relative Chow stability.
@@
This is an equality whenever $\mathcal{F}$ is finitely generated because, in that case, the approximation $\mathcal{F}_{(r)}$ coincides with $\mathcal{F}$, viewed as a filtration of $(X,rL)$. Moreover, as $\mathcal{F}^{\lambda}_{(r)}S_{r}=\mathcal{F}^{\lambda}S_{r}$ for any filtration $\mathcal{F}$, the inner product satisfies $\langle\mathcal{F}_{(r)},\beta_{i}\rangle_{1}=\langle\mathcal{F},\beta_{i}\rangle_{r}$. Putting everything together proves the result. ∎
Assume now that $p$ is fixed by $T$. We obtain a torus $\widehat{T}$ in $\mathrm{Aut}(\widehat{Y},\widehat{H}_{m})$ which leaves $\widehat{\mathcal{F}}$ invariant. Combining Theorem 3.4 and Proposition 3.5, we obtain the following relative formula:
Theorem 3.8.
Assume that $G_{\mathcal{F}}\leq\Lambda$ on $\Delta_{\frac{1}{m}}$ as in Theorem 3.4. Then,
$\mathrm{Chow}_{r,T}(\widehat{\mathcal{F}})\leq\mathrm{Chow}_{1,T}(\mathcal{F}_{(mr)})+\Big(\nu_{T}(p,\mathcal{F})-\frac{b_{0}}{a_{0}}\Big)\frac{m^{1-n}}{2(n-2)!a_{0}}+o(m^{-n}r^{0})$
@@
(3.4) $\nu_{T}(p,\mathcal{F})=\nu(p,\mathcal{F})-\sum_{i=1}^{d}\frac{\langle\mathcal{F},\beta_{i}\rangle}{\langle\beta_{i},\beta_{i}\rangle}\Big(\nu(p,\beta_{i})-\frac{b_{i,0}}{a_{0}}\Big),$
given any $\mathbb{Z}$-basis $(\beta_{i})$ of $N_{\mathbb{Z}}(T)$ and $w_{\beta_{i}}(k)=b_{i,0}k^{n+1}+O(k^{n})$.
@@
In order to use the formula in Proposition 3.8 for an instability argument on the blowup, we need to find an invariant point $p$ for which the relative weight in Equation (3.4) is small enough for large $m$. For this, we establish a $T$-invariant version of results of Boucksom [Szé15, Section 8].
Let $S^{\prime}$ be a graded subalgebra of the ring of sections $S=\bigoplus_{k\geq 0}S_{k}$ for $S_{k}=H^{0}(Y,kH)$, not equal to $\mathbb{C}$. Recall the volume of $S^{\prime}$ is the limit
@@
where $\mathfrak{bs}(S_{k}^{\prime})$ is the base-ideal of the linear series $S_{k}^{\prime}$ in $S$. The proofs of Lemmas 3.11 and 3.12 follow verbatim the proofs of [Szé15, Lemma 22] and [Szé15, Lemma 21] noting that the divisorial valuations and their centers constructed in the proofs are invariant. We therefore omit the details.
@@
Let $\Lambda<e_{\mathrm{max}}(\mathcal{F})$ be a real number. There exists a $T$-invariant point $p$ in $X$ such that for any Okounkov body $P$ centered at $p$, there is an integer $m_{1}\geq m_{0}$ satisfying
$G_{\mathcal{F}}|_{\Delta_{\frac{1}{m_{1}}}}\leq\Lambda.$